OpenStudy (anonymous):

I want to build a right triangular garden on the side of my house. Find the sides of the triangle if the hypotenuse is 6 feet and the two sides are equal in length.

OpenStudy (anonymous):

 Pythagorean Theorem: a² + b² = c² \[x² + x² = 6²\] \[2x² = 36\] \[x² = 18\] \[x = \sqrt{18}\] \[x=sqrt{9}*sqrt{2}\] \[x = 3sqrt{2}\]

OpenStudy (anonymous):

There is a much simpler method to this, but this is down to the fundamentals.